OpenStudy (anonymous):

So, h(t) is twice-differentiable with these values: When t=0, h(t)=1 and h'(t)=-9 t=1, h(t)=-4 and h'(t)=1 t=2, h(t)=-2 and h'(t)=3 t=3, h(t)=-1 and h'(t)=1 t=4, h(t)=-4 and h'(t)=-8 Justifying your reasoning, answer the following: a) Determine where h(t) = 0 on the interval (0, 4). b) Determine where h'(t) = 0 on the interval (0, 4). c) Determine where h"(t) = 0 on the interval (0,4).

OpenStudy (anonymous):

where as in "exactly for what value of t" or where as in "in what interval"?

OpenStudy (anonymous):

for example, by the mean value theorem we know that there exists a c in the inteval (1,3) for which \[h''(c)=\frac{h'(3)-h'(1)}{3-1}=\frac{1-1}{2}=0\] but that just tells us that h'' as a zero in the interval (1,3) not exactly where it is. unless i miss my guess you cannot know exactly for what value of c is h''(c) = 0

OpenStudy (anonymous):

maybe i am missing something

OpenStudy (anonymous):

That's what I thought, I can find the intervals but not the exact value and I thought I was just going crazy. But I dont know what they mean to ask

OpenStudy (anonymous):

really unless i am totally clueless i cannot imagine that it is asking for exact values

OpenStudy (anonymous):

so i am going to assume the "where" means "on what interval in (0,4)" not "for what number"

OpenStudy (anonymous):

in which case it should be clear by mvt right?

OpenStudy (anonymous):

Yeah. If its an interval, I think I got it. Thanks!